OpenStudy (anonymous):

. Solve the absolute value equation. Check your solutions. | 5x + 13| = –7

OpenStudy (khaakenajaf):

not possible

OpenStudy (anonymous):

not possible

OpenStudy (anonymous):



OpenStudy (khaakenajaf):

because when the absolute bracets open they make the other side +-.. here the other side is already with - sign

OpenStudy (anonymous):

show the work

OpenStudy (anonymous):

what work? it's impossible. the absolute value brackets indicate a number a certain disctance away from 0, so the number can not be a negative number

OpenStudy (anonymous):

5x+13=-7 5x=20 x=-4

OpenStudy (anonymous):

that's not the correct answer because you're ignoring the absolute value

OpenStudy (anonymous):

5x+13=-7 5x=-7-13 5x=-20 divide both sides by 5 x=-4

OpenStudy (anonymous):

I understand the algebra, but you are ignoring the absolute value sign. That makes it an impossible equation. You can not get a negative answer from an equation inside an absolute value.

OpenStudy (anonymous):

5x+13 maybe positive, zero, or negative. :5x+13>=0 and 5x+13<=0 if 5x+13>=0 then {5x+13}=5x+13-7 5x=-7-13 5x=-20 x=-4

OpenStudy (anonymous):

If 5x+13<0 then |5x+13|=-(5x+13)=-7 -5x-13=-7 5x+13=7 5x=7-13 5x=-6 x=-6/5 ANSWER: x is either -4 or -6/5

OpenStudy (khaakenajaf):

listen @sharna u are not understanding any of us when the constant already bears a negative sign then the working or solution is not possible

OpenStudy (anonymous):

it can't be negative when it's in absolute value brackets though. It will always be a positive distance from zero. I think you need to look up what absolute value is

OpenStudy (anonymous):

thx. just trying to figure it out at all angles